Circular No. 8395

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 Q2 (MACHHOLZ) J. J. Gonzalez, Leon, Spain, reports that a visual observation made

from Alto del Castro with a 0.20-m reflector on Aug. 28.16 UT yielded total mag 10.9 and coma diameter 2′.3. P. Birtwhistle, Great Shefford, Berkshire, England, reports that CCD images obtained with a 0.30-m f/6.3 Schmidt-Cassegrain reflector on Aug. 28.15 show a coma of diameter 70′′ and a broad tail 5′ long in p.a. 240^{o}. R. Behrend, Geneva Observatory, reports that CCD images by C. Vuissoz and himself with the 1.20-m Euler reflector on Aug. 28.4 show a coma diameter of 1′ and a 3′ tail in p.a. 240^{o}.

Additional astrometry and the following preliminary parabolic orbital elements appear on MPEC 2004-Q58 (ephemeris: H = 6.0, n = 7.5).

T = 2005 Jan. 27.370 TT  ω = 14.^{o}807
Ω = 95.921 }2000.0
q = 1.28614 AU  i = 38.326
2004 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Aug. 23 4^{h}10.^{m}35 −21^{o}59.^{′}7 2.268 2.545 93.^{o}9 23.^{o}3 10.8
28 4 16.91 −22 24.6 2.174 2.492  96.0 23.8 10.7
Sept. 2 4 23.31 −22 52.1 2.082 2.438  98.1 24.2 10.5
7 4 29.51 −23 22.3 1.991 2.385 100.2 24.6 10.3
12 4 35.48 −23 55.0 1.901 2.331 102.2 25.0 10.2
17 4 41.18 −24 29.8 1.813 2.278 104.2 25.3 10.0
22 4 46.56 −25 06.3 1.727 2.225 106.1 25.7  9.8
27 4 51.58 −25 44.2 1.642 2.172 108.0 26.0  9.6
Oct. 2 4 56.20 −26 22.9 1.559 2.120 109.9 26.4  9.4

SUPERNOVAE 2004dy AND 2004dz M. Ganeshalingam, M. Moore, and W. Li report the LOSS (cf. IAUC

8388) discoveries of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004dy Aug. 25.25 21^{h}11^{m}30.^{s}13 − 2^{o}02′13′′.3 19.4 5′′.8 W, 16′′.8 S
2004dz Aug. 20.52 2 57 36.66 −10 10 52.8 18.6 0′′.3 E, 2′′.9 N

Additional KAIT magnitudes: SN 2004dy in IC 5090, Aug. 21.24 UT, [20.0; 27.26, 18.8. SN 2004dz, 2003 Oct. 29.43, [20.0; Aug. 27.55, 18.6.

2004 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green